OpenStudy (studygurl14):

HELP??!! Tin reacts with fluorine to form two different compounds, A and B. Compound A contains 38.5 g of tin for each 12.3 g of fluorine. Compound B contains 56.5 g of tin for each 36.2 g of fluorine. What is the lowest whole-number mass ratio of tin that combines with a given mass of fluorine?

OpenStudy (perl):

compound A T38.5 F 12.3 divide both by 12.3 T (38.5/12.3) F( 12.3/12.3)

OpenStudy (perl):

That gives approximately T3 F

OpenStudy (perl):

the next one T (56.5 / 36.2 ) F (36.2 / 36.2) T(1.56 ) F 1 but you don't want decimals so T ( 2*1.5 ) F (2*1) T 3 F2

OpenStudy (studygurl14):

How'd you get from 1.56 to 3 again?

OpenStudy (surry99):

lowest whole number...so multiply by two

OpenStudy (here_to_help15):

Compound.......Mass of Sn........Mass of F ....... A ................. 38.5 g .............. 12.3 g ....... B ................. 56.5 g .............. 36.3 g We must fix the masses of flourine in both compounds. For this you can; - either take the mass of fluorine 12.3 g in both compounds - either take the mass of fluorine 36.3 g in both compounds - or you can the mass of fluorine any arbitrary value in both compounds. But in the last choice you have to calculate the masses of Sn in both compounds. However, in the first two choices, you have to find only the single mass of Sn. Suppose let's chose the first one. If 56.5 g Sn combines with 36.3 g F in compound B, What will be the mass of Sn combining with 12.3 g F ? Mass of Sn = 56.5 x (12.3 / 36.3) = 19.14 g Sn. Now we have; Compound.......Mass of Sn........Mass of F ....... A ................. 38.5 g .............. 12.3 g ....... B ................. 19.14 g .............12.3 g Mass ratio of tin that combines with a fixed mass of fluorine: 38.5 / 19.14 = 2 You can also try the other alternatives. They will always give 2.

OpenStudy (here_to_help15):

I may be bad in French but never bad in Math

OpenStudy (perl):

1.56 ~ 1.5 , and we want whole numbers .

OpenStudy (perl):

1.5 = 3/2 , so multiply by 2

OpenStudy (studygurl14):

So..it would be 2:1 tin to fluorine?

OpenStudy (perl):

you have to do it to both

OpenStudy (perl):

T(1.56 ) F1 2*( T(1.56 ) F1) = T(2*1.5)F(1*2) = T3 F2

OpenStudy (here_to_help15):

Yes @perl is correct you must do both, and if you ever need more help in Chemistry just send me a message me :D

OpenStudy (here_to_help15):

Take me off at the end lol put it twice by accident

OpenStudy (studygurl14):

I'm a little confused still...sorry

OpenStudy (here_to_help15):

Hm lemme see if can explain it in a different way

OpenStudy (studygurl14):

I get that it has to be a whole number, but why multiply by 2? can't you just round?

OpenStudy (here_to_help15):

I am putting as snap shot becuz it wont let me put terms

OpenStudy (studygurl14):

ok

OpenStudy (studygurl14):

where'd you get the 118... and the 18.99...?

OpenStudy (here_to_help15):

-.- lol im getting a headache lemme check

OpenStudy (studygurl14):

lol srry